Compute the largest eigenvalue in magnitude, and the corresponding eigenvector of
the matrix
A=1 −1 1
2 0 3
1 4 −1
by performing four iterations of the power method
1
Expert's answer
2020-04-22T19:30:49-0400

"A=\begin{bmatrix}
 1&-1&1 \\
 2&0&3\\
1&4&-1
\end{bmatrix}"

Begin with an initial nonzero approximation of

"x_0=\begin{bmatrix} 1 \\ 1\\ 1 \end{bmatrix}"

Then obtain the following approximations

Iteration

"x_1=Ax_0=\begin{bmatrix} 1 & -1&1 \\ 2 & 0&3\\ 1&4&-1 \end{bmatrix}\begin{bmatrix} 1 \\ 1
\\ 1 \end{bmatrix} =\begin{bmatrix} 1 \\ 5\\ 4 \end{bmatrix}"

“Scaled” Approximation

"\to1\begin{bmatrix} 1 \\ 5\\ 4 \end{bmatrix}"

"x_2=Ax_1=\begin{bmatrix} 1 & -1&1 \\ 2 & 0&3\\ 1&4&-1 \end{bmatrix}\begin{bmatrix} 1 \\ 5\\ 4 \end{bmatrix} =\begin{bmatrix} 0 \\ 14\\ 17 \end{bmatrix}"

"\to14\begin{bmatrix} 0 \\ 1\\ 1.2 \end{bmatrix}"

"x_3=Ax_2=\begin{bmatrix} 1 & -1&1 \\ 2 & 0&3\\ 1&4&-1 \end{bmatrix}\begin{bmatrix} 0 \\ 14\\ 17 \end{bmatrix} =\begin{bmatrix} 3 \\ 51\\ 39 \end{bmatrix}\\
\to3\begin{bmatrix} 1 \\ 17\\ 13 \end{bmatrix}\\
x_4=Ax_3=\begin{bmatrix} 1 & -1&1 \\ 2 & 0&3\\ 1&4&-1 \end{bmatrix}\begin{bmatrix} 3 \\ 51\\ 39 \end{bmatrix} =\begin{bmatrix} -9 \\ 123\\ 168 \end{bmatrix}\\
\to-9\begin{bmatrix} 1 \\ -13.7\\ -18.7 \end{bmatrix}"

Note that the approximations appear to be approaching scalar multiples of

"\begin{bmatrix} 1 \\ -14\\ -19 \end{bmatrix}"

With x=(1,-13.7,-18.7) as the approximation of a dominant eigenvector of A , 

use the Rayleigh quotient to obtain an approximation of the dominant eigenvalue of A . First compute the product Ax

"Ax=\begin{bmatrix} 1 & -1&1 \\ 2 & 0&3\\ 1&4&-1 \end{bmatrix}\begin{bmatrix} 1 \\ -13.7\\ -18.7 \end{bmatrix} =\begin{bmatrix} -4 \\ -54.1\\ -35.1 \end{bmatrix}"

Then, because

"Ax\cdot x=-4\cdot 1+(-54.1)(-13.7)+(-35.1)(-18.7)=1393.54\\
 x\cdot x=1\cdot1+(-13.7)(-13.7)+(-18.7)(-18.7)=538.38"

you can compute the Rayleigh quotient to be

"\lambda=\frac{Ax\cdot x}{x\cdot x}=\frac{1393.54}{538.38}=2.6"

which is a good approximation of the dominant eigenvalue

"\lambda=3"

Answer:

"\begin{bmatrix} 1 \\ -14\\ -19 \end{bmatrix}, \lambda=3"